# Compact and Selective Disclosure for Verifiable Credentials

### v1

@@ Abstract
Self-Sovereign Identity (SSI) is a novel identity model that empowers individuals with full control over their data, enabling them to choose what information to disclose, with whom, and when. This paradigm is rapidly gaining traction worldwide, supported by numerous initiatives such as the European Digital Identity (EUDI) Regulation or Singapore’s National Digital Identity (NDI). For instance, by 2026, the EUDI Regulation will enable all European citizens to seamlessly access services across Europe using Verifiable Credentials (VCs). A key feature of SSI is the ability to selectively disclose only specific claims within a credential, enhancing privacy protection of the identity owner. This paper proposes a novel mechanism designed to achieve Compact and Selective Disclosure for VCs (CSD-JWT). Our method leverages a cryptographic accumulator to encode claims within a credential to a unique, compact representation. We implemented CSD-JWT as an open-source solution and extensively evaluated its performance under various conditions. CSD-JWT provides significant memory savings, reducing usage by up to 46% compared to the state-of-the-art. It also minimizes network overhead by producing remarkably smaller Verifiable Presentations (VPs), reduced in size by 27% to 93%. Such features make CSD-JWT especially well-suited for resource-constrained devices, including hardware wallets designed for managing credentials.
Index Terms:
Self-Sovereign Identity (SSI), Verifiable Credentials (VCs), Selective Disclosure, Digital Identity, Privacy.
@@ I Introduction
The growing emphasis on preserving user privacy has driven a significant transformation in the digital identity ecosystem, shifting from centralized to decentralized models [1]. Emerging privacy-preserving regulations further reinforce this evolution [2, 3, 4] that place individuals at the center of the identity management process. This paradigm shift has given rise to Self-Sovereign Identity (SSI) [5], a novel identity model that is gaining worldwide traction. In February 2024, the European Commission approved the ”electronic IDentification, Authentication, and trust Services” (eIDAS 2.0) [6], mandating the creation of the European Digital Identity Wallet (EUDI Wallet) [7]. Under this framework, by 2026, European member states must provide citizens with a digital identity wallet for managing their credentials. Similarly, Singapore’s National Digital Identity (NDI) system [8], implemented through Singpass, enables citizens to seamlessly access public and private services while maintaining full ownership of personal data.
@@
The core principle of SSI consists of empowering individuals with full control over their data by allowing them to decide what information to disclose, with whom, and when, eliminating reliance on centralized authorities. This is achieved through Verifiable Credentials (VCs) [9], digital credentials that attest specific properties or attributes of the holder, hereinafter called claims. Unlike traditional Public Key Infrastructure (PKI) certificates [10], which provide secure public key distribution and rely on centralized authorities, VCs are designed for broader use cases and operate within a decentralized framework. These credentials are cryptographically signed by trusted issuers such as governments, enabling decentralized verification of the stored information. A key feature of SSI is the ability for identity owners to selectively disclose only the specific subset of claims contained in VCs. For instance, individuals can prove a legal drinking age without revealing unnecessary information like their address.
@@
Various VC formats have been proposed [11], with JSON Web Token (JWT) [12] emerging as one of the most widely adopted as it offers a concise approach to exchanging verifiable claims. To enable privacy-preserving selective disclosure, the Internet Engineering Task Force (IETF) proposed the Selective Disclosure for JWTs (SD-JWT) [13] specification. SD-JWT replaces plain text claims with digests of their salted values. The issuer signs the resulting JWT and shares it with the holder, along with the original claims and salts. To disclose data, the holder provides the verifier with the signed JWT and salt-claim pairs corresponding to disclosed claims.
@@
Problem Definition. A major downside of this method is that credential size grows linearly with the number of claims, increasing end-user storage and network requirements. Moreover, although digests hide undisclosed information, SD-JWT reveals the exact number of included claims, raising privacy concerns. This knowledge can potentially expose the identity owner to inference attacks [14, 15] as the number of claims may be exploited to infer details about hidden information.
@@
These drawbacks of SD-JWT increase when VCs are applied to entities beyond individuals, such as Internet of Things (IoT) devices [16, 17, 18]. Given the heterogeneous capabilities of these devices, considerations around storage, computation, and networking are even more crucial. In constrained environments, resources are shared across different functions. For example, memory must store both VCs and firmware, making even KB-level reductions impactful. Moreover, minimizing resource usage is essential for enhancing energy efficiency, a critical factor in resource-constrained settings. These concerns also apply to hardware wallets, known for their enhanced security in managing VCs [19, 20], which often have limited storage capacities ranging from a few KB to a couple of MB [21]. This highlights the need for further selective disclosure mechanisms that minimize on-device storage requirements and optimize data transmission.
@@
CSD-JWT. This paper proposes a mechanism that achieves Compact and Selective Disclosure for VCs (CSD-JWT), diminishing storage and network overhead of SD-JWT. Our method gives identity owners full control over their information while minimizing the amount of data to store and share. CSD-JWT replaces the full list of salted hashes with a fixed-length value, generated by accumulating all claims into a cryptographic accumulator [22]. This ensures that the credential size remains constant regardless of the number of claims it contains. Furthermore, it hides the exact number of claims in the VC, offering stronger resilience against inference attacks.
@@
The issuer must provide proof of inclusion for each claim to be disclosed, known as witness. The holder combines these witnesses with corresponding claims to generate Verifiable Presentations (VPs). The storage requirements for disclosing claims, i.e., CSD-JWT and witnesses, are lower than those for SD-JWT and salts due to the constant size of our credentials. This property also reduces the amount of data shared with the verifier, making CSD-JWT an efficient solution for constrained environments.
@@
Evaluation. We implemented a prototype of CSD-JWT and conducted a series of experiments, including tests on an RFC7228 class-2 constrained device [23], to compare our method with the state-of-the-art mechanism, i.e., SD-JWT. In particular, we measured storage and computational requirement to evaluate CSD-JWT suitability for constrained devices. It is worth noting that our evaluation considers a holder maintains a single credential. However, as underscored by the EUDI Regulation [4], users are likely to use multiple VCs by 2026 in real-world scenarios [24], making the overhead reduction offered by CSD-JWT even more impactful. Experimental findings proved that our approach can achieve up to 46% memory savings for each stored credential. Experiments on the network overhead shows that VP in CSD-JWT is 93% smaller when the holder aims to disclose only 1 claim out of 100 (maximum privacy), while it still provides a 27% reduction when no privacy preservation is required and all claims are revealed.
Contributions. Our work has the following contributions:
@@
Organization. The remainder of this paper is structured as follows: Section II provides the background on VCs and cryptographic accumulators. Section III presents the reference system and threat model. Section IV details how CSD-JWT achieves compact and selective disclosure for VCs. Section V analyzes the security of the proposed mechanism in addressing the identified threats. Section VI comprehensively evaluates CSD-JWT. Section VII reviews literature in the field. Finally, Section VIII draws our conclusions.
@@ II Background
II-A Verifiable Credential (VC)
A digital identity can be envisioned as a set of attributes or properties, which provides a snapshot of any entity, including individuals, companies, and objects. Any third party must be able to directly verify these attributes through attestations, embedded in digital credentials issued by trusted authorities. A data structure that satisfies this property is referred to as a VC. Several VC data models have been proposed in recent years, among which the World Wide Web Consortium (W3C) VCDM [25] and SD-JWT VC [26] being considered the most prominent. We propose CSD-JWT as a selective disclosure mechanism that can be seamlessly integrated into any model, offering a practical alternative to SD-JWT.
To be effective and address the challenges of centralized identity models, a verifier (e.g., a service provider) should be able to directly verify the authenticity of the claims contained, without direct interaction with the issuer. This necessitates that the issuer identity can be directly retrieved from the credential itself. Consequently, every party must have a globally unique self-generated digital identity [27]. The commonly adopted approaches are public keys and Decentralized Identifiers (DIDs) [28]. A DID is a Uniform Resource Identifier (URI) standardized by the W3C. Each DID is associated with a DID Document, encompassing publicly available information (e.g., public key) about the entity it identifies. The DID Document is typically shared in DID-based scenarios through verifiable data registries such as a blockchain [29]. Digital identifiers are crucial for identifying issuers and holders, preventing the reuse of previously collected VCs, which may allow an adversary to impersonate the identity owner. In CSD-JWT, digital identification for both holders and issuers is achieved through DIDs. However, it could be easily extended to support public key identification.
@@ II-B Cryptographic Accumulators
Cryptographic accumulators are cryptographic primitives that securely map a set of values in a fixed-size output called accumulator value [30]. Each accumulated element can be associated with a proof of inclusion, known as a witness. This enables verifying the membership without disclosing any other elements of the set. Although several implementations have been proposed, all types of accumulators must satisfy two properties:
@@
We apply $h_{l}$ to all the $e_{i}$ to be included in the accumulator, obtaining the accumulator value $a$. For each $e_{i}$ there exists a witness $w_{i}$ such that $a=h_{l}(e_{i},w_{i})$, proving its inclusion within $a$. The one-way property ensures that given the accumulator value $a$, it is computationally infeasible to find a $w_{i}$ that proves membership for an $e_{i}$ that does not belong to the set. Moreover, using a function that meets the quasi-commutative requirement guarantees that the accumulator remains independent of the order in which values are accumulated.
Accumulators can be classified based on their support for verifying element inclusion. Specifically, those that enable verification using membership witnesses are called positive, as opposed to those supporting non-membership witnesses which are known as negative. The ones that support both types of verification methods are called universal. Various accumulators have been proposed over the years [31, 32, 33, 34, 35]. Among these, accumulators based on Bilinear Groups [32, 33, 22] are widely acknowledged for their efficiency [36, 37], especially in terms of witness size and verification efficiency for inclusion. Given these properties, CSD-JWT leverages a positive ECC-based accumulator to hide claims within VCs.
@@ II-C Selective Disclosure
Selective disclosure is a mechanism that allows an identity owner $h$ to disclose only a subset of claims $C\subseteq\mathcal{C}$ from a verifiable credential to a verifier $v$.
@@ III-A Reference System
Issuer. An issuer is a trusted entity such as a government, university, or municipality that issues VCs to identity owners. For example, an issuer may provide an individual with a VC serving as a driving license. The issuer signs credentials with its private key, making it directly verifiable to third parties. It is worth noting that a crucial assumption of the SSI model is that holders and verifiers trust the issuer.
@@ III-B Threat Model
The security of CSD-JWT depends on the secure generation of the accumulator value and witnesses. As the issuer is a trusted entity for all the parties involved in an SSI ecosystem, we can safely assume that they have been securely computed, ensuring resistance to collisions. In this subsection, we provide an overview of the adversary perspective and outline the potential threats to our system, based on the threat model presented in [9].
@@
Adversary’s Capability. We adopt the Dolev-Yao adversary model [38], which assumes that the adversary can eavesdrop on, intercept, and inject an unlimited number of messages.
@@
  • Replay Attacks: An adversary may present a valid credential that belongs to another identity owner. This VC could be obtained through collusion with another holder or stolen from an unsuspecting identity owner. A service provider may wrongly grant authorization believing claims, which do not represent the adversary.
  • Data Overcollection: A service provider may collect more claims than necessary to provide a service, either with malicious intent or unintentionally. Moreover, the service provider may be able to infer additional details on the identity owner given side information such as the number of claims in the VC.
@@ IV-A Preliminaries
Accumulator. We aggregate all claims $\mathcal{C}$, within a verifiable credential $vc$ belonging to an identity owner $h$, into a single value $a$. The elements included in $a$ are derived by applying a hash function $h()$ to each $c_{j}\in\mathcal{C}$. Specifically, we use the cryptographically secure SHA-256 hash function, which is resistant to collisions [39]. CSD-JWT employs the positive ECC-based accumulator presented in [22], which supports batch operations for accumulating elements and generating witnesses. The accumulator value is mapped as a 256-bit point over the BN254 curve.
@@
$vc$ Verifiable credential
$vp$ Verifiable presentation
@@
Functions. In the following, we report the functions used to realize CSD-JWT:
  • $par\leftarrow\mathsf{Setup}(1^{\lambda})$: Given the security parameter $1^{\lambda}$, which determines the level of security, this function initializes and outputs the accumulator parameters $par$. Accumulator parameters comprise the random generator of the ECC curve $\mathcal{G}$, and a key pair ($pk_{i},sk_{i})$.
  • $a\leftarrow\mathsf{AccumulatorSetup}(par)$: This function receives accumulator parameters $par$ and setups the accumulator, generating the initial accumulator value $a$.
  • $a\leftarrow\mathsf{AccumulateClaims}(a,\mathcal{C},sk_{i})$: This function takes the set of claims $\mathcal{C}$, the private key $sk_{i}$, and maps $c_{j}\in\mathcal{C}$ onto the accumulator elliptic curve, aggregating them into a single accumulator value $a$.
  • $\mathcal{W}\leftarrow\mathsf{ComputeWitnesses}(a,\mathcal{C},sk_{i})$: Given the accumulator value $a$, the set of claims $C$, the private key $sk_{i}$, this function generates the set $\mathcal{W}$, where each $w_{j}\in\mathcal{W}$ corresponds to a $c_{j}\in\mathcal{C}$.
  • $0,1\leftarrow\mathsf{VerifyWitness}(a,c_{j},w_{j},pk_{i})$: This function verifies whether the claim $c_{j}$ was accumulated in $a$. This verification requires both the witness $w_{j}$ corresponding to $c_{j}$ and the accumulator public key $pk_{i}$.
@@
  • $0,1\leftarrow\mathsf{VerifyPresentation}(vp)$: The verifier retrieves the public keys of $pk_{h}$ and $pk_{i}$ through their respective DID using a $\mathsf{Resolve()}$ function, which allows obtaining corresponding DID Documents. The verifier checks the presentation ownership through $\mathsf{Verify(vp,pk_{h})}$. Finally by with the $\mathsf{VerifyWitness}()$ function, the verifier confirms whether the provided claims are included in the accumulator value $a$ contained in $vp$.

### v2

# Alessandro Buldini1, Carlo Mazzocca2, Rebecca Montanari1, Selcuk Uluagac3

@@ Abstract
Self-Sovereign Identity (SSI) is a novel identity model that empowers individuals with full control over their data, enabling them to choose what information to disclose, with whom, and when. This paradigm is rapidly gaining traction worldwide, supported by numerous initiatives such as the European Digital Identity (EUDI) Regulation or Singapore’s National Digital Identity (NDI). For instance, by 2026, the EUDI Regulation will enable all European citizens to seamlessly access services across Europe using Verifiable Credentials (VCs). A key feature of SSI is the ability to selectively disclose only specific claims within a credential, enhancing the privacy protection of the identity owner. This paper proposes a novel mechanism designed to achieve Compact and Selective Disclosure for VCs (CSD-JWT). Our method leverages a cryptographic accumulator to encode claims within a credential into a unique, compact representation. We implemented CSD-JWT as an open-source solution and extensively evaluated its performance under various conditions. CSD-JWT provides significant memory savings, lowering usage by up to 46% compared to the state-of-the-art. It also minimizes network overhead by producing remarkably smaller Verifiable Presentations (VPs), with size reduction from 27% to 93%. Such features make CSD-JWT especially well-suited for resource-constrained devices, including hardware wallets designed for managing credentials.
@@ I Introduction
The growing emphasis on preserving user privacy has driven a significant transformation in the digital identity ecosystem, shifting from centralized to decentralized models [mazzocca2024survey]. Emerging privacy-preserving regulations further reinforce this evolution [gdpr2016, californiaccpa, eu2024regulation1183] that place individuals at the center of the identity management process. This paradigm shift has given rise to Self-Sovereign Identity (SSI) [MUHLE201880], a novel identity model that is gaining worldwide traction. In February 2024, the European Commission approved the ”electronic IDentification, Authentication, and trust Services” (eIDAS 2.0) [eidas2.0], mandating the creation of the European Digital Identity Wallet (EUDI Wallet) [wallet]. Under this framework, by 2026, European member states must provide citizens with a digital identity wallet for managing their credentials. Similarly, Singapore’s National Digital Identity (NDI) system [singaporeNDI], implemented through Singpass, enables citizens to seamlessly access public and private services while maintaining full ownership of personal data.
@@
The core principle of SSI consists of empowering individuals with full control over their data by allowing them to decide what information to disclose, with whom, and when, eliminating reliance on centralized authorities. This is achieved through Verifiable Credentials (VCs) [krul2024sok], digital credentials that attest specific properties or attributes of the holder, hereinafter called claims. Unlike traditional Public Key Infrastructure (PKI) certificates [pki_survey], which provide secure public key distribution and rely on centralized authorities, VCs are designed for broader use cases and operate within a decentralized framework. These credentials are cryptographically signed by trusted issuers (e.g., governments), enabling decentralized verification of the stored information. A key feature of SSI is the ability for identity owners to selectively disclose only a specific subset of claims contained in VCs. For instance, individuals can prove a legal drinking age without revealing unnecessary information like their address.
@@
Various VC formats have been proposed [format_vc], with JSON Web Token (JWT) [jones2015rfc] emerging as one of the most widely adopted, as it offers a concise approach to exchanging verifiable claims. To enable privacy-preserving selective disclosure, the Internet Engineering Task Force (IETF) proposed the Selective Disclosure for JWTs (SD-JWT) [SD-JWT] specification. SD-JWT replaces plain text claims with digests of their salted values. The issuer signs the resulting JWT and shares it with the holder, along with the original claims and salts. To disclose data, the holder provides the verifier with the signed JWT and salt-claim pairs corresponding to disclosed claims.
@@
Problem Definition. A major downside of this method is that credential size grows linearly with the number of claims, increasing end-user storage and network requirements. Moreover, although digests hide undisclosed information, SD-JWT reveals the exact number of included claims, raising privacy concerns. This knowledge can potentially expose the identity owner to inference attacks [mehnaz2022your, 8068648] as the number of claims may be exploited to infer details about hidden information.
@@
These drawbacks become especially critical when VCs are applied to entities beyond individuals, such as Internet of Things (IoT) devices [mazzoccaevoke, fotiu, mahalle2020rethinking]. Given the heterogeneous capabilities of these devices, considerations around storage, computation, and networking are even more crucial. In constrained environments, resources are limited and shared across different functions. For example, memory must store both VCs and firmware, making even KB-level reductions impactful. Moreover, minimizing resource usage is essential for enhancing energy efficiency, a critical factor in resource-constrained settings. These concerns also apply to hardware wallets, known for their enhanced security in managing VCs [ccshardware, 9122595], which often have limited storage capacities ranging from a few KB to a couple of MB [ledgermemory]. This raises the research question: How can selective disclosure mechanisms balance computation, storage, network overhead, and privacy in resource-constrained environments?
@@
CSD-JWT. This paper proposes a protocol that achieves Compact and Selective Disclosure for VCs (CSD-JWT), diminishing storage and network overhead of SD-JWT. Our method gives identity owners full control over their information while minimizing the amount of data to store and share. CSD-JWT replaces the full list of salted hashes with a fixed-length value, obtained by aggregating all claims into a cryptographic trapdoor-based accumulator [benaloh_demare]. This significantly reduces the credential size: rather than storing a separate hash and salt for each claim, CSD-JWT requires only a single accumulator value for the entire credential and one witness per claim, i.e., a proof of inclusion.
It is worth noting that the design of CSD-JWT is agnostic of the specific accumulator implementation. This flexibility allows the underlying accumulator to be replaced as more secure, efficient, or space-optimized constructions emerge. The only essential requirement for the accumulator employed in CSD-JWT is that it must be trapdoor-based. Furthermore, with the substitution of the signed list of hashes with a single accumulator value, CSD-JWT allows to hide the exact number of claims in the VC, offering stronger resilience against inference attacks. In contrast, SD-JWT requires the use of decoy digests to achieve similar privacy guarantees, which further increases the overall size of the credential [ietf-oauth-selective-disclosure-jwt-22].
@@
In CSD-JWT, the issuer must provide a witness for each claim to be disclosed. The holder then combines these witnesses with corresponding claims to generate Verifiable Presentations (VPs). The storage requirements for disclosing claims, i.e., accumulator value and witnesses, are lower than those for signed list and salts due to the constant size of the accumulator value. This property also reduces the amount of data shared with the verifier, making CSD-JWT an efficient solution for constrained environments.
@@
Evaluation. We implemented a prototype of CSD-JWT based on the cryptographic accumulator proposed in [vitto_biryukov], and conducted a series of experiments, including tests on an RFC7228 class-2 constrained device [rfc7228], to compare our method with the state-of-the-art mechanism, i.e., SD-JWT. In particular, we measured storage and computational requirement to evaluate CSD-JWT suitability for constrained devices. It is worth noting that our evaluation considers a holder who maintains a single credential. However, as underscored by the EUDI Regulation [eu2024regulation1183], users are likely to use multiple VCs by 2026 in real-world scenarios [gartner2024digitalidentity], making the overhead reduction offered by CSD-JWT even more impactful. Experimental findings proved that our approach can achieve up to 46% memory savings for each stored credential. Experiments on the network overhead shows that VP in CSD-JWT is 93% smaller when the holder aims to disclose only 1 claim out of 100 (maximum privacy), while it still provides a 27% reduction when no privacy preservation is required and all claims are revealed.
Contributions. In the following, we summarize the key contributions of this work:
@@
Organization. The remainder of this paper is structured as follows. Section II provides the background on VCs and cryptographic accumulators. Section III presents the reference system and threat model. Section IV details how CSD-JWT achieves compact and selective disclosure for VCs. Section V analyzes the security of the proposed mechanism in addressing the identified threats. Section VI comprehensively evaluates CSD-JWT. Section VII reviews literature in the field. Finally, Section VIII draws our conclusions.
@@ II Background
II-A Verifiable Credential
A digital identity can be envisioned as a set of attributes or properties that provides a snapshot of any entity, including individuals, companies, and objects. Any third party must be able to directly verify these attributes through attestations, embedded within digital credentials issued by trusted authorities. A data structure that satisfies this property is referred to as a VC. Several VC data models have been proposed in recent years, among which the World Wide Web Consortium (W3C) VCDM [w3c_dm] and SD-JWT VC [sd_jwt_vc] are considered the most prominent. We propose CSD-JWT as a selective disclosure mechanism that can be seamlessly integrated into any model, offering a practical alternative to SD-JWT.
To be effective and address the challenges of centralized identity models, a verifier (e.g., a service provider) should be able to directly verify the authenticity of the claims contained, without direct interaction with the issuer. This necessitates that the issuer identity can be directly retrieved from the credential itself. Consequently, every party must have a globally unique self-generated digital identity [sokdatasover]. The commonly adopted approaches are public keys and Decentralized Identifiers (DIDs) [did]. A DID is a Uniform Resource Identifier (URI) standardized by the W3C. Each DID is associated with a DID Document, encompassing publicly available information (e.g., public key) about the entity it identifies. The DID Document is typically shared through verifiable data registries such as a blockchain [10.1145/3465481.3469204]. Digital identifiers are crucial for identifying issuers and holders, preventing the reuse of previously collected VCs, which may allow an adversary to impersonate the identity owner. In CSD-JWT, digital identification for both holders and issuers is achieved through DIDs. However, it could be easily extended to support public key identification.
@@ II-B Cryptographic Accumulators
Cryptographic accumulators are cryptographic primitives that securely map a set of values in a fixed-size output called the accumulator value [benaloh_demare]. Each accumulated element can be associated with a proof of inclusion, known as a witness. This enables verifying the membership without disclosing any other elements of the set. Although several implementations have been proposed, all types of accumulators must satisfy two properties:
@@
We apply $h_{l}$ to all the elements $e_{i}$ to be included in the accumulator, obtaining the accumulator value $a$. For each $e_{i}$ there exists a witness $w_{i}$ such that $a=h_{l}(e_{i},w_{i})$, proving its inclusion within $a$. The one-way property ensures that given the accumulator value $a$, it is computationally infeasible to find a $w_{i}$ that proves membership for an $e_{i}$ that does not belong to the set. Moreover, using a function that meets the quasi-commutative requirement guarantees that the accumulator remains independent of the order in which values are accumulated.
Accumulators can be classified based on their support for verifying element inclusion. Specifically, those that enable verification using membership witnesses are called positive, as opposed to those supporting non-membership witnesses, which are known as negative. The ones that support both types of verification methods are called universal. Various accumulators have been proposed over the years [camenish_lysyanskaya, nguyen, au, papamanthou, boneh_merkle_trees]. Among these, accumulators based on Bilinear Groups [nguyen, au, vitto_biryukov] are widely acknowledged for their efficiency [wangetal, kumaretal], especially in terms of witness size and verification efficiency for inclusion. Given these properties, in Section VI, we evaluate CSD-JWT by instantiating it with a positive ECC-based accumulator.
@@ II-C Selective Disclosure
Selective disclosure is a mechanism that allows an identity owner $h$ to disclose only a subset of claims $C\subseteq\mathcal{C}$ to a verifier $v$ so that $C$ is still verifiable.
@@ III-A Reference System
Issuer. An issuer is a trusted entity, such as a government, university, or municipality, that issues VCs to identity owners. For example, an issuer may provide an individual with a VC serving as a driving license. The issuer signs credentials with its private key, making it directly verifiable to third parties. It is worth noting that a crucial assumption of the SSI model is that holders and verifiers trust the issuer.
@@ III-B Threat Model
In SSI, issuers are typically assumed to be trusted by all parties involved [krul2024sok], as they certify claims (of which they are inherently aware) that verifiers rely on to grant access to resources and services. Accordingly, in CSD-JWT, the generation of the accumulator value and corresponding witnesses is securely managed by a trusted issuer. Our threat model aligns with that of SD-JWT, which does not guarantee unlinkability in scenarios where trusted issuers collude with other entities [baum2024cryptographers]. This limitation stems from the fact that hashes, salts, accumulator values, and witnesses remain constant across credential presentations, making repeated disclosures linkable under adversarial conditions. A more in-depth analysis of unlinkability and its implications can be found in Appendix A. In this subsection, we provide an overview of the adversary perspective and outline the potential threats to our system, based on the threat model presented in [krul2024sok].
@@
Adversary’s Capability. We adopt the Dolev-Yao adversary model [dolevyao], which assumes that the adversary can eavesdrop on, intercept, and inject an unlimited number of messages.
@@
  • Replay Attacks: An adversary may present a valid credential that belongs to another identity owner. This VC could be obtained through collusion with another holder or stolen from an unsuspecting identity owner. A service provider may wrongly grant authorization, believing claims that do not represent the adversary.
  • Data Overcollection: A service provider may collect more claims than necessary to provide a service, either with malicious intent or unintentionally. Moreover, the service provider may be able to infer additional details on the identity owner, given side information such as the number of claims in the VC.
@@ IV-A Preliminaries
Accumulator. We aggregate all claims $\mathcal{C}$, within a verifiable credential $vc$ belonging to an identity owner $h$, into a single value $a$. The elements included in $a$ are derived by applying a hash function $h()$ to each $c_{j}\in\mathcal{C}$. Specifically, we use the cryptographically secure SHA-256 hash function, which is resistant to collisions [sha256].
@@
vcvc Verifiable credential
vpvp Verifiable presentation
@@
Functions. In the following, we report the functions used to realize CSD-JWT. All functions related to setup, as well as accumulator and witness generation, are performed by the issuer.
  • $par\leftarrow\mathsf{Setup}(1^{\lambda})$: Given the security parameter $1^{\lambda}$, which determines the level of security, this function initializes and outputs the accumulator parameters parpar that comprises a key pair ($pk_{i},sk_{i})$ and all the necessary parameters required by the underlying cryptographic accumulator instance.
  • $a\leftarrow\mathsf{AccumulatorSetup}(par)$: This function receives accumulator parameters parpar and setups the accumulator, generating the initial accumulator value $a$.
  • $a\leftarrow\mathsf{AccumulateClaims}(a,\mathcal{C},sk_{i})$: This function takes the set of claims $\mathcal{C}$, the private key $sk_{i}$, and aggregates the claims $c_{j}\in\mathcal{C}$ into the accumulator value $a$.
  • $\mathcal{W}\leftarrow\mathsf{ComputeWitnesses}(a,\mathcal{C},sk_{i})$: Given the accumulator value $a$, the set of claims $\mathcal{C}$, the private key $sk_{i}$, this function generates the set $\mathcal{W}$, where each $w_{j}\in\mathcal{W}$ corresponds to a $c_{j}\in\mathcal{C}$.
  • $0,1\leftarrow\mathsf{VerifyWitness}(a,c_{j},w_{j},pk_{i})$: The verifier $v$ checks whether the claim $c_{j}$ was accumulated in $a$. This verification requires both the witness $w_{j}$ corresponding to $c_{j}$ and the accumulator public key $pk_{i}$.
@@
  • $0,1\leftarrow\mathsf{VerifyPresentation}(vp)$: The verifier $v$ retrieves the public keys of $pk_{h}$ and $pk_{i}$ through their respective DID using a $\mathsf{Resolve()}$ function, which allows obtaining corresponding DID Documents. The verifier checks the presentation ownership through $\mathsf{Verify(vp,pk_{h})}$. Then, it extracts the accumulator value $a$ and all claim-witness pairs $(c_{j},w_{j})$ from the $vp$ before calling the $\mathsf{VerifyWitness}()$ function to confirm whether the provided claims are included in $a$.